# ICE-SEARCH: A Language Model-Driven Feature Selection Approach

### v2

@@ Abstract
This study unveils the In-Context Evolutionary Search (ICE-SEARCH) method, the first work that melds language models (LMs) with evolutionary algorithms for feature selection (FS) tasks and demonstrates its effectiveness in Medical Predictive Analytics (MPA) applications. ICE-SEARCH harnesses the crossover and mutation capabilities inherent in LMs within an evolutionary framework, significantly improving FS through the model’s comprehensive world knowledge and its adaptability to a variety of roles. Our evaluation of this methodology spans three crucial MPA tasks: stroke, cardiovascular disease, and diabetes, where ICE-SEARCH outperforms traditional FS methods in pinpointing essential features for medical applications. ICE-SEARCH achieves State-of-the-Art (SOTA) performance in stroke prediction and diabetes prediction; the Decision-Randomized ICE-SEARCH ranks as SOTA in cardiovascular disease prediction. Our results not only demonstrate the efficacy of ICE-SEARCH in medical FS but also underscore the versatility, efficiency, and scalability of integrating LMs in FS tasks. The study emphasizes the critical role of incorporating domain-specific insights, illustrating ICE-SEARCH’s robustness, generalizability, and swift convergence. This opens avenues for further research into comprehensive and intricate FS landscapes, marking a significant stride in the application of artificial intelligence in medical predictive analytics.
@@ 1 Introduction
The proposed In-Context Evolutionary Search (ICE-SEARCH) method is the first work that melds anguage models (LMs) with evolutionary algorithms for feature selection (FS) tasks and demonstrates its effectiveness in Medical Predictive Analytics (MPA) applications.
Medical Predictive Analytics (MPA) represents a critical and specialized sector within health informatics dedicated to the sophisticated examination of medical data. MPA leverages advanced statistical algorithms and machine learning methodologies to forecast upcoming trends in healthcare outcomes, including patient admission rates, pharmaceutical advancements, and strategies for disease management (Van Calster et al. (2019)).
The core value of MPA resides in its ability to sift through and analyze the extensive datasets prevalent in the healthcare industry. This analytical power supports more informed clinical decision-making, enhances patient care quality, and facilitates more efficient allocation of healthcare resources. (Lamba et al. (2021)).
However, the sheer volume and complexity of this data pose significant challenges in converting it into actionable insights. In this context, the process of feature selection (FS) becomes critically important. By identifying the most pertinent subset of features within vast datasets, FS facilitates a more efficient analysis. Specifically, in medical scenarios like stroke prediction, where data is both extensive and varied, FS proves invaluable. It not only circumvents the "curse of dimensionality" but also aids in pinpointing disease causes. This, in turn, can unveil a new understanding of disease mechanisms, thereby contributing to advancements in medical research and patient care strategies (Remeseiro and Bolón-Canedo (2019)).
As highlighted in Section 2, FS confronts substantial hurdles, chiefly due to its broad search space that leads to an exponential increase in computational demands. Complications such as diminished generalizability, redundancy within intricate problem scenarios, challenges in handling high-dimensional data, an incomplete understanding of underlying data structures, and the obligation to presume data distributions that may not hold true, further exacerbate these challenges.
Contrasting this, the utility of Language Models (LMs) has significantly evolved, extending far beyond traditional text analysis. This evolution positions LMs as critical instruments in specialized fields. The adaptability and advanced capabilities of these models are demonstrated by their leading-edge performance in various sectors, including medical diagnostics Singhal et al. (2022), financial analysis Wu et al. (2023), data preprocessing Borisov et al. (2023), code generation Rozière et al. (2024) etc. This broad applicability underscores the potential of LMs to revolutionize not only how we approach complex datasets but also how we extract and interpret the most meaningful insights from them.
We posit that LMs can effectively contribute to FS in various applications. Furthermore, as suggested by Chen et al. (2023), semantic knowledge encoded in the LM allows for access to a considerable amount of cross-domain knowledge that would not be available in evolutionary algorithms, greater diversity of the outputs, and reduces the amount of manual design and human bias involved in the evolutionary algorithm, thereby powering evolutionary neuro-architecture search.
Hence, we introduce ICE-SEARCH, a FS methodology that integrates the in-context learning capabilities of language models Dong et al. (2023) with an evolutionary algorithmic framework. Our method involves engaging the LM in diverse roles, asking it to identify crucial features for a specific downstream task, and providing the LM with feedback on the performance of both effective and ineffective feature sets, measured in terms of training and validation accuracies of 10-fold cross-validation(CV). We also discover that during evolutionary in-context training, there is an enhancement in the model’s ability to uncover a set of high-quality feature combinations. At the time of submitting this manuscript, our research appears to be the first to explore the potential of LMs in this context.
To demonstrate the efficacy of our approach, we analyze the performance of ICE-SEARCH on three critical tasks–stroke prediction, cardiovascular disease prediction, and diabetes prediction–using three increasingly complex datasets. Each of these diseases is among the leading causes of morbidity and mortality worldwide, carrying substantial risks of severe health complications, long-term disability, and death.
We demonstrate that ICE-SEARCH consistently outperforms the existing FS methods in stroke prediction and diabetes prediction, and can be easily adapted to improve its performance on cardiovscular disease prediction whose dataset suffers from a discrepancy between training/valitaion and testing distribution. More interestingly, our empirical studies show that ICE-SEARCH is robust against poor population initialization and converges rapidly.
@@
    ICE-SEARCH is the first work to apply an LM to dynamically search for valuable feature sets through role-playing and improve its in-context prompting examples without any fine-tuning.
@@
    ICE-SEARCH achieves SOTA, our experiments on stroke prediction and diabetes prediction demonstrate the effectiveness, robustness, stability, and scalability over existing approaches.
@@
    The value of ICE-SEARCH lies in its ability to empower end-users with cross-domain expertise encoded in a LM fine-tuned with domain-specific knowledge. ICE-SEARCH can easily be adapted to perform FS in areas like genetic analysis that commonly have thousands or even millions of features.
@@ 2 Related Work
Feature Selection: FS presents itself as a difficult task mainly due to two challenges: (i) its vast search space, where the total possible solutions are $2^{n}$ for a dataset with n features, leading to exponential growth in computing expenses, and (ii) the sub-optimal convergence caused by stagnations in local optima. Most of the techniques applied, such as heuristic search and random search, suffer from poor generalizability of evaluation criteria, redundant feature subsets when applied to complex problems Xue et al. (2016a), huge search space for high-dimensional data whose efficiency deteriorates quickly or computationally infeasible Miao and Niu (2016); Li et al. (2017), limited knowledge of the underlying data strucutre Li et al. (2017), or imposing case-specific assumption(s) about the probability of the distribution of the data Miao and Niu (2016). Despite that evolutionary computation(EC) methods have shown promises Xue et al. (2016a), flexible evaluation measures according to specific feature representations and search mechanisms that incorporate worldly knowledge across multiple disciplines are still in need for their optimal performance Li et al. (2017). Feature selection algorithms that are scalable, robust against dataset quality, and efficient are preferred Doquet and Sebag (2020); Li et al. (2017).
Evolutionary Algorithms for Feature Selection: Due to the vast search space, finding an optimal feature subset remains a challenging task. As discussed in the survey written by Chandrashekar and Sahin (2014), various ML and statistical FS algorithms have been proposed to tackle the task. In particular, as suggested by Xue et al. (2016b), EC-based approaches have recently been proven to be effective in many scenarios. Specifically, EC methods for feature selection have been applied in scenarios including genetic algorithm Holland (1992), particle swarm optimization Wang et al. (2018), genetic programming Muni et al. (2006), and correlation-based filter measures Hall (1999). Apart from the prospect of evolutionary algorithms for FS, it points out several issues and challenges that need to be addressed: scalability, computation cost, search mechanisms, stability, etc. Specifically, the FS algorithms need to: (1) have the capacity to process thousands or even millions of features, (2) have an efficient search strategy and a fast evaluation measure, (3) explore the whole search space and exploit the local regions if needed, and (4) be stable when using different starting points.
LLM for Feature Engineering: Feature engineering is closely related to feature selection, the former focuses on constructing new features from the original feature pool while the latter aims to identify and remove redundant and informative features. The article by Hollmann et al. (2023) introduces Context-Aware Automated Feature Engineering (CAAFE), a method leveraging LLMs for automated feature engineering in tabular datasets. CAAFE uses LLMs to iteratively generate both the Python codes for creating new semantically meaningful features and explanations for their utility. The method improved performance on 11 out of 14 datasets, demonstrating its effectiveness over traditional methods. Similarly, CAAFE also combines the domain knowledge encapsulated in LLMs with classical machine learning algorithms, offering a novel approach to integrating domain knowledge into automated machine learning processes.
Evolutionary Prompting: Prompting is a technique employed to give LLMs leverage to address specialized tasks more efficiently. However, the performance of LLMs varies contingent upon the selection of the prompt, and that leads to the topic of automatic prompt optimizations, such as continuous prompt tuning and discrete prompts Liu et al. (2023); Li and Liang (2021); Zhang et al. (2021). In contrast, the high variance in prompt effectiveness for downstream tasks has led to the development of prompt generation, prompt revision, and prompt editing methods that focus on the exploration or exploitation of candidate prompts to select the most optimal one(s) Zhou et al. (2022); Guo et al. (2023a); Zhang et al. (2022). To address the constraints of prompts designed for predefined answers, which may encounter challenges when applied to generation tasks, and to mitigate the risk of encountering local optima, recently introduced evolutionary prompting techniques offer the advantages of not relying on predefined parameters or gradients while simultaneously maintaining a balance between exploration and exploitation Guo et al. (2023b); Chen et al. (2023). Such paradigms of evolutionary prompting inspired ICE-Search, which adapts and evolves in the search for optimal prompts to identify critical features for downstream tasks. Similar to evolutionary prompting, ICE-Search showcases novel ways of leveraging language models while offering flexibility and adaptability to diverse metrics and tasks.
@@ 3 ICE-SEARCH: Feature Selection with Language Model
As evidenced by section 4.5, the identified features are indicative of the risk factors of diseases and can improve the accuracy of MPA. ICE-Search’s ability to extract representative features from medical datasets emphasizes its crucial role in MPA. This capability allows healthcare professionals to develop precise and efficient predictive models, enabling the optimization of healthcare resources for early disease detection. In this section we explain our algorithm in detail.
Notations
  • $\mathcal{X}$: A mapping that uses a set of X number of classical FS methods used to initialize the feature set pool. Output X feature sets.
  • Y: Number of times for the LM to select feature sets through zero-shot prompting for feature set pool initializaion.
  • $\mathcal{Z}$: Set of role-plays for the LM intended for evolution.
  • U: Number of top feature sets to keep.
  • V: Number of bottom feature sets to keep.
  • $N$: Number of groups that the given dataset is split into for CV.
  • $E$: Number of Epochs.
  • $\mathcal{F}$: The given feature set
  • $\mathcal{S}$: Feature set pool (i.e. population)
  • $\mathcal{D}$: Dataset.
  • $\mathcal{T}$: Task description.
  • $\Phi(\mathcal{D},s,N)$: Evaluation of the input feature set $f$ usig a pre-determined downstream model on dataset $\mathcal{D}$, output average training and testing accuracies across N-fold CV.
[…]
Figure 1: An illustration of ICE-SEARCH Architecture
@@ 3.1 Algorithm Archetecture
ICE-SEARCH architecture can be broken down into four components: Initialization, Crossover and Mutation, Filtration, Final Selection. Figure 1 provides a high-level overview of our algorithm and Algorithm 1 illustrates its low-level implementation.
Initialization (line 4-12): The first stage of ICE-SEARCH is the initialization of a population $\mathcal{S}$, which is a pool of feature sets. This initial population is derived using $\mathit{X}$ number of classical FS methods, with each method’s performance assessed based on average training and validation accuracies of the downstream model $\Phi$ over N-fold CV. To further refine the initial population, we prompt a LM through $\Phi(\cdot)$ $Y$ times in a zero-shot manner, and the output feature combination that yields the highest average validation accuracy over N-fold CV is chosen and put in $\mathcal{S}$ for further processing.
Crossover and Mutation (line 13-20): In view of the achievements attributed to EVO Prompting by Chen et al. (2023) and EVO Prompt by Guo et al. (2023b), our evolutionary strategy leverages a LM as the mutation and crossover operator throughout each epoch. Specifically, the LM assumes $\mathcal{Z}$ distinct roles and is informed about the average training and validation accuracies associated with each feature set in $\mathcal{S}$. During each of these role-plays $\Psi$, the LM selects a set of features, and its average training and validation accuracies over N-fold CV is computed via $\Phi(\cdot)$ and recorded.
@@
Final Selection (line 24-28): The iterative evolutionary process is repeated for E epochs. Upon completion of the final epoch, we select the feature set that demonstrates the highest validation accuracy. In cases where multiple feature combinations exhibit the same highest validation accuracy, the tie-breaker is the combination with the lowest training accuracy. This approach ensures that our model not only performs well on unseen data but also avoids overfitting to the training dataset. On the other hand, as shown in table 5 and 8, we have found empirically that all $\mathit{U}$ feature sets are valuable as many of them yield high test accuracy. Therefore, it is not mandatory to follow strictly the strategy we just described. One can customize this component by adjusting step 24 to 27 in algorithm 1 according to the underlying task complexity or dataset quality.

### v4

@@ Abstract
This study unveils the In-Context Evolutionary Search (ICE-SEARCH) method, which is among the first works that melds large language models (LLMs) with evolutionary algorithms for feature selection (FS) tasks and demonstrates its effectiveness in Medical Predictive Analytics (MPA) applications. ICE-SEARCH harnesses the crossover and mutation capabilities inherent in LLMs within an evolutionary framework, significantly improving FS through the model’s comprehensive world knowledge and its adaptability to a variety of roles. Our evaluation of this methodology spans three crucial MPA tasks: stroke, cardiovascular disease, and diabetes, where ICE-SEARCH outperforms traditional FS methods in pinpointing essential features for medical applications. ICE-SEARCH achieves State-of-the-Art (SOTA) performance in stroke prediction and diabetes prediction; the Decision-Randomized ICE-SEARCH ranks as SOTA in cardiovascular disease prediction. The study emphasizes the critical role of incorporating domain-specific insights, illustrating ICE-SEARCH’s robustness, generalizability, and convergence. This opens avenues for further research into comprehensive and intricate FS landscapes, marking a significant stride in the application of artificial intelligence in medical predictive analytics.
@@ 1 Introduction
The proposed In-Context Evolutionary Search (ICE-SEARCH) method is among the first works that meld large language models (LLMs) with evolutionary algorithms for feature selection (FS) tasks and demonstrates its effectiveness in Medical Predictive Analytics (MPA) applications.
As a data pre-processing technique, FS provides efficient and effective data preparation for data mining and machine learning problems (Li et al. (2017a)). This procedure can reduce the number of necessary features and facilitate accuracy prediction (Jain and Zongker (1997); Reddy and Kumar (2016)). For example, FS proves invaluable in stroke prediction, where data is extensive and varied as they are crucial for identifying key predictors among a large number of stroke risk factors (Capela et al. (2015); Zhang et al. (2018); Fang et al. (2020)). Feature selection (FS) is challenging due to its vast search space, escalating to $2^{n}$ possible solutions for n features, which exponentially increases computing costs. Common issues with FS techniques include handling large, high-dimensional data inefficiently, limited understanding of data structure, and assumptions about data distribution that may not apply universally. Thus, the evolutionary computation(EC)-based approach (or evolutionary algorithms (EAs)) becomes one of the popular FS algorithms, as it has proven to be an effective solution for feature selection given its global search ability( Xue et al. (2016)). The EC-based approach treats each FC result as a potential candidate and improves these candidates by manually-designed crossover rules.
The use of Large Language Models (LLMs) has expanded beyond traditional text analysis (Zhao et al. (2023)), demonstrating versatility and high performance across various sectors, including medical diagnostics ( Singhal et al. (2022)), financial analysis ( Wu et al. (2023)), data preprocessing ( Borisov et al. (2023)), and code generation ( Rozière et al. (2024)). This broad applicability highlights LLMs’ potential to transform our approach to complex datasets. Specifically, various research has shown the great benefit of using LLM as a cross-over operator in EAs Chen et al. (2023). As previous work ( Guo et al. (2023a)) suggests, EAs typically modify tokens independently in sequences and ignore the essential connections critical for coherence. In this method, LLMs generate new prompts while EAs oversee the optimization process to select the optimal prompts. This approach reduces the amount of manual design and human bias involved in the evolutionary algorithm, thereby powering, for instance, evolutionary neuro-architecture search (Chen et al. (2023)). Moreover, semantic knowledge encoded in the LLM allows for access to a considerable amount of cross-domain knowledge (Ye et al. (2021); Sanh et al. (2021); Weller et al. (2020)) that would not be available in evolutionary algorithms. More importantly, such knowledge can be enhanced through role-play prompting (Kong et al. (2024)).
Hence, based on the success of EC-based FS approaches and LLM as the cross-over operator, we introduce ICE-SEARCH, an FS methodology that integrates the in-context learning (Brown et al. (2020)) and role-playing capabilities of language models with an evolutionary algorithmic framework. Our method involves engaging the LLM in diverse roles, asking it to identify crucial features for a specific downstream task, and providing the LLM with feedback on the performance of both effective and ineffective feature sets via in-context few-shot prompting, measured in terms of training and validation accuracies of 10-fold cross-validation(CV). Our method improves on traditional feature selection algorithms by merging evolutionary computation with pre-trained knowledge from a large language model. This integration helps us identify important features with improved global search capabilities and validate them without assuming data distribution or requiring a comprehensive understanding of the data structure. To demonstrate the efficacy of our approach, we analyze ICE-SEARCH’s performance on the tasks of MPA: stroke prediction, cardiovascular disease prediction, and diabetes prediction- using three increasingly complex datasets.
@@
    ICE-SEARCH is the first work to apply an LLM to iteratively search for valuable feature sets through role-playing and improve its in-context prompting examples without fine-tuning.
@@
    By integrating an LLM’s inherent domain knowledge with an evolutionary algorithm, ICE-SEARCH effectively reduces the search space in MPA datasets and guides the random crossover procedure in an EC-based feature selection approach through its domain knowledge.
@@
    ICE-SEARCH achieves SOTA performance in feature selection. Our experiments on three widely adopted MPA tasks demonstrate its effectiveness and robustness.
@@ 2 Related Work
Feature Selection and its Evolutionary Solutions: Feature Selection (FS) is a critical dimensionality reduction technique used in data mining and machine learning by selecting a subset of significant features for model construction( Zebari et al. (2020)), thereby preserving the original meanings of these features ( Chandrashekar and Sahin (2014a)). Its primary goal is to mitigate the impact of noise or irrelevant features in datasets, ensuring the accuracy and efficiency of predictions in subsequent models ( Zebari et al. (2020)). In practice, FS presents itself as a difficult task mainly due to its vast search space, where the total possible solutions are $2^{n}$ for a dataset with n features, leading to exponential growth in computing expenses( Lyu et al. (2023)). For example, most of the techniques applied suffer from: (i) huge search space for high-dimensional data whose efficiency deteriorates quickly or computationally infeasible( Miao and Niu (2016); Li et al. (2017b)), (ii) limited knowledge of the underlying data structure( Li et al. (2017b)), and (iii) imposed case-specific assumption(s) about the probability of the distribution of the data( Miao and Niu (2016)). To solve (i), evolutionary computation (EC)-based approaches are developed (Chandrashekar and Sahin (2014b)) as they are able to effectively search the space of possible feature subsets (Kim et al. (2000)). Specifically, EC methods for feature selection have been applied in scenarios including genetic algorithmHolland (1992)), particle swarm optimizationWang et al. (2018)), genetic programmingMuni et al. (2006)), and correlation-based filter measuresHall (1999)). On the other hand, LLMs have shown great knowledge generalizability (Wei et al. (2021)) and led to significant research work such as LLM-augmented knowledge graph (Pan et al. (2024)). Thus, our approach surpasses traditional EC methods by combining EC optimization techniques with pre-trained domain knowledge from an LLM. This integration allows us to leverage the pre-trained knowledge to identify significant feature subsets, which we then validate through a downstream task. Unlike other methods, ours does not presuppose any assumptions about the underlying data distribution, effectively addressing issues (i) through (iii).
Prompting and Evolutionary Prompting: Prompting is a technique employed to give LLMs leverage to address specialized tasks more efficiently. Various state-of-the-art prompting techniques, such as zero-shot or few-shot in-context learning (ICL)(Brown et al. (2020); Min et al. (2021); Dong et al. (2023a, b)), role-play prompting, (Xu et al. (2023); Zheng et al. (2023); Kong et al. (2024)), Chain-of-Thought(COT) (Wei et al. (2022)), Tree-of-Thought(TOT) (Yao et al. (2023a)), and Graph-of-Thought(GOT) (Besta et al. (2023)) have been proven effective in tackling various tasks such as cross-lingual multi-step reasoning Ranaldi and Zanzotto (2023).
However, the performance of LLMs varies depending upon the selection of the prompt ( Liu et al. (2023)), and that leads to automatic prompt optimizations, such as continuous prompt tuning ( Li and Liang (2021)) and differentiable prompts( Zhang et al. (2021)). In contrast, the high variance in prompt effectiveness for downstream tasks has led to the development of prompt engineering ( Zhou et al. (2022)), prompt revision ( Guo et al. (2023b)), and prompt editing ( Zhang et al. (2022)) methods that focus on the exploration or exploitation of candidate prompts to select the most optimal one(s).
To address the constraints of prompts designed for predefined answers and mitigate the risk of encountering local optima, many have explored evolutionary prompting optimization techniques ( Yang et al. (2023a); Meyerson et al. (2023); Guo et al. (2023a); Chen et al. (2023)). They offer the advantages of not relying on predefined parameters or gradients while simultaneously maintaining a balance between exploration and exploitation.
Such paradigms of evolutionary prompting inspired ICE-Search, which adapts and evolves in the search for optimal prompts via role-playing different experts to identify critical features for downstream tasks. Similar to evolutionary prompting, ICE-Search showcases novel ways of leveraging language models via ICL while offering flexibility and adaptability to diverse metrics and tasks.
Medical Predictive Analysis: Medical Predictive Analytics (MPA) represents a critical and specialized sector within health informatics dedicated to sophisticated medical data examination. MPA leverages advanced statistical algorithms and machine learning methodologies to forecast upcoming trends in healthcare outcomes, including patient admission rates, pharmaceutical advancements, and strategies for disease management (Van Calster et al. (2019)). For example, previous works have focused on leveraging various algorithms to facilitate predictive data mining for medical diagnosis( Soni et al. (2011)), disease prediction( Usama et al. (2019)), robotic surgery( Oniani et al. (2020)), use of disease diagnosis systems for Electronic Medical Records( Latif et al. (2020)), as well as analysis and prediction of outpatient visits( Jin et al. (2019)).
3 Preliminary
In this section, we formulate the feature selection problem. FS is a process of choosing a subset of the original features based on an evaluation criterion that assesses the optimality of each subset. The selection process can be generalized into four steps( Liu and Yu (2005)), as illustrated in Figure 1: subset generation, subset evaluation, stopping criterion, and result validation.
Figure 1: Feature Selection 4-step Process (adapted from Liu and Yu (2005))
Consider a target task $T$ and a dataset $D$ composed of input-output pairs $(x,y)\in D$ relevant to task $T$. Define a feature space $F$ and a probability distribution $\pi_{\theta}:F\rightarrow\{0,1\}$ parameterized by $\theta$, from which we can sample feature subsets $S\subseteq F$. Additionally, we define an evaluation function $\Phi(S,D):2^{F}\times D\rightarrow\mathbb{R}$ that applies a learning algorithm to train a model using the feature subset $S$ on $D$, subsequently outputting a real-valued score $s$. This score $s$ reflects the effectiveness of $S$ in terms of accuracy and other performance metrics. The primary objective is to discover a set of feature subsets $S\sim 2^{F}$ that, when used to train models on $D$, optimize the evaluation metric $\Phi(S,D)$.
The feature selection process includes four key steps:
  1. 1.
    Subset Generation: Using the probability distribution $\pi_{\theta}$, this step can employ various search strategies to generate candidate feature subsets $S\subseteq F$.
  2. 2.
    Subset Evaluation: Each generated subset $S$ is evaluated based on a predefined evaluation criterion $\Phi(S,D):2^{F}\times D\rightarrow\mathbb{R}$, which assesses how well the model, trained using features in subset $S$ on dataset $D$, performs with respect to the target task $T$. The criterion can include measures of model accuracy, complexity, etc.
  3. 3.
    Stopping Criterion: A stopping criterion indicates a sufficient level of performance and terminates the process of subset generation, and evaluation is repeated. For example, the margin of improvement between iterations, as calculated using a predefined evaluation metric, serves as a stopping criterion.
  4. 4.
    Result Validation: The best-performing feature subset $S^{*}$ is validated and further tested on unseen data or using different validation techniques to ensure the subset’s generalizability and prone to overfitting.
Overall, FS identifies an optimal subset $S^{*}$ of features from $F$ that, when used to train a model on $D$, maximizes the evaluation function $\Phi(S,D)$ and thus enhances the model’s performance on task $T$.
@@ 4 ICE-SEARCH: Feature Selection with Language Model
In this section, we explain our algorithm in detail.
@@ 4.1 Algorithm Archetecture
ICE-SEARCH architecture can be broken down into four components: Initialization, Crossover and Mutation, Filtration, Final Selection. Figure 2 provides a high-level overview of our algorithm and Algorithm 1 illustrates its low-level implementation. Notation definitions can be found in Appendix A.1
Figure 2: An illustration of ICE-SEARCH Architecture
Initialization: The first stage of ICE-SEARCH is the initialization of a population $\mathcal{S}$, which is a pool of feature sets. This initial population is derived using $\mathit{X}$ number of classical FS methods, with each method’s performance assessed based on average training and validation accuracies of the downstream model $\Phi$ over N-fold CV. To further refine the initial population, we prompt an LLM $Y$ times in a zero-shot manner, and the output feature combination that yields the highest average validation accuracy over N-fold CV is chosen and put in $\mathcal{S}$ for further processing. This is discussed in (line 4-12).
Crossover and Mutation (line 13-20): In view of the achievements attributed to EVO Prompting by Chen et al. (2023) and EVO Prompt by Guo et al. (2023a), our evolutionary strategy leverages an LLM as the mutation and crossover operator throughout each epoch. Specifically, the LLM assumes $\mathcal{Z}$ distinct roles and is informed about the average training and validation accuracies associated with each feature set in $\mathcal{S}$. During each of these role-plays $\Psi$, the LLM selects a set of features, and its average training and validation accuracies over N-fold CV is computed via $\Phi(\cdot)$ and recorded.
@@
Final Selection (line 24-28): The iterative, evolutionary process is repeated for E epochs. Upon completing the final epoch, we select the feature set demonstrating the highest validation accuracy. In cases where multiple feature combinations exhibit the same highest validation accuracy, the tie-breaker is the combination with the lowest training accuracy. This approach ensures that our model performs well on unseen data and avoids overfitting the training dataset. On the other hand, as shown in Table 5 and 8, we have found empirically that all $\mathit{U}$ feature sets are valuable as many of them yield high test accuracy. Therefore, it is not mandatory to follow strictly the strategy we just described. One can customize this component by adjusting steps 24 to 27 in Algorithm 1 according to the underlying task complexity or dataset quality.